# Lie Neurons: Adjoint-Equivariant Neural Networks for Semisimple Lie Algebras

### v2

@@ Abstract
This paper proposes an equivariant neural network that takes data in any semi-simple Lie algebra as input. The corresponding group acts on the Lie algebra as adjoint operations, making our proposed network adjoint-equivariant. Our framework generalizes the Vector Neurons, a simple $\mathrm{SO}(3)$-equivariant network, from 3-D Euclidean space to Lie algebra spaces, building upon the invariance property of the Killing form. Furthermore, we propose novel Lie bracket layers and geometric channel mixing layers that extend the modeling capacity. Experiments are conducted for the $\mathfrak{so}(3)$ and $\mathfrak{sl}(3)$ Lie algebras on various tasks, including fitting equivariant and invariant functions, learning system dynamics, point cloud registration, and homography-based shape classification. Our proposed equivariant network shows wide applicability and competitive performance in various domains.
@@ 1 Introduction
The equivariance property preserves the symmetry group structure, often a Lie group, such that the feature map commutes with the group representation. Equivariant models have gained success in various domains, including but not limited to the modeling of molecules (Thomas et al., 2018), physical systems (Finzi et al., 2020), social networks (Maron et al., 2018), images (Worrall et al., 2017), and point clouds (Zhu et al., 2023b). Convolutional neural networks (CNNs) are translation-equivariant, enabling stable image features regardless of the pixel positions in the image plane. Typical extensions include rotation (Cohen et al., 2017) and scale (Worrall & Welling, 2019) equivariance, while more general extensions are also explored (MacDonald et al., 2022).
Figure 1: The Lie Neurons can be applied in learning dynamics. In this example, we learn the dynamics of a simulated free-rotation International Space Station, which we pose as an initial value problem in the Neural ODE framework. This figure shows the estimated trajectories and the learned vector field from the Lie Neurons. Detail descriptions can be found in Section 5.1.2.
In this paper, we propose a new type of equivariant model that captures the symmetry in Lie algebra spaces. For a given Lie algebra, we represent its elements as vectors by specifying a set of bases, and represent the adjoint actions as matrix multiplications, exploiting the isometry structure of Lie algebras to handle such data in typical vector form. Through the connection between inner products and the Killing form, we generalize the architecture of Vector Neurons (Deng et al., 2021), a $\mathrm{SO}(3)$-equivariant network originally designed for point cloud data in 3-D Euclidean space, to process data in arbitrary semisimple Lie algebra. We further build new types of equivariant network layers that exploit the structure of Lie algebras and extend the flexibility of the model. Using $\mathfrak{so}(3)$ and $\mathfrak{sl}(3)$ as examples, we conduct experiments on various tasks in physics, computer vision, and general function fitting, and show that our proposed network has wide applicability and competitive performance.
@@
    We propose a new adjoint-equivariant network architecture, enabling the processing of semisimple Lie algebraic input data. Such models frequently appear in real-world geometric problems.
@@
    The software implementation can be found at https://github.com/UMich-CURLY/LieNeurons.
@@ 2 Related Work
Equivariant networks enable the model output to change in a predicted way as the input goes through certain transformations. This means that the model, by construction, generalizes over the variations caused by those transformations. Therefore, it reduces the sampling complexity in learning and improves the robustness and transparency facing input variations. (Cohen & Welling, 2016a) initials the idea to generalize equivariance in deep learning models, realizing equivariance to 90-degree rotations in a 2D image plane using group convolution. This method works with discrete transformations by augmenting the input domain with a dimension for the set of transformations. The approach is generalized to other discretized groups in $SE(2)$, $SE(3)$, and $E(3)$ (Hoogeboom et al., 2018; Winkels & Cohen, 2018; Worrall & Brostow, 2018; Chen et al., 2021; Zhu et al., 2023a). Steerable convolution is proposed in (Cohen & Welling, 2016b), leveraging the irreducible representations to remove the need for discretization and facilitate equivariant convolution on continuous groups in the frequency domain (Worrall et al., 2017; Cohen et al., 2017; Weiler et al., 2018). Beyond convolutions, more general equivariant network architectures are proposed, for example, (Fuchs et al., 2020; Hutchinson et al., 2021; Chatzipantazis et al., 2022) for transformers and (Batzner et al., 2022; Brandstetter et al., 2021) for message passing networks. Vector Neurons (Deng et al., 2021) designs a multi-layer perception (MLP) and graph network that generalize the scalar features to 3D features to realize $\mathrm{SO}(3)$-equivariance on spatial data. The above works mainly focus on compact groups, on which more general recipes for building equivariant layers that are not limited to a specific group are also proposed (Kondor & Trivedi, 2018; Cohen et al., 2019; Weiler & Cesa, 2019; Xu et al., 2022; Lang & Weiler, 2020). The extension of equivariance beyond compact groups is also explored. (Finzi et al., 2021) constructs MLPs equivariant to arbitrary matrix groups using their finite-dimensional representations. With the Monte Carlo estimator, equivariant convolutions are generalized to matrix groups with surjective exponential maps (Finzi et al., 2020) and all finite-dimensional Lie groups (MacDonald et al., 2022), where Lie algebra is used to parameterize elements in the continuous Lie groups as a lifted domain from the input space.
Our model structure resembles the MLP style of Vector Neurons (Deng et al., 2021), but our work models the equivariance of arbitrary semisimple groups under adjoint actions. Different from Vector Neurons, the input domain of our method is the Lie algebra.
@@ 3 Preliminaries
We provide some preliminaries for Lie groups by focusing on matrix Lie groups. For detailed explanations, we refer the readers to (Hall, 2013; Rossmann, 2006; Kirillov, 2008).
@@ 3.1 Lie Group and Lie Algebra
A Lie group $\mathcal{G}$ is a smooth manifold whose elements satisfy the group axioms. Because of this, a special vector space naturally arises at the identity of every Lie group named the Lie algebra, denoted $\mathfrak{g}$. A Lie algebra locally captures the structure of the Lie group.
Every Lie algebra is equipped with an asymmetric binary operator called the Lie bracket:
@@
The elements of the Lie algebra have non-trivial structures. However, since the Lie algebra is a vector space, for a Lie algebra of dimension $m$, we can always represent it using $\mathbb{R}^{m}$ given appropriate bases. As a result, we introduce two useful maps (Chirikjian, 2011):
@@
where $e_{i}$ are the canonical basis of $\mathbb{R}^{m}$ and $E_{i}=(e_{i})^{\wedge}\in\mathfrak{g}$.
@@ 3.2 Adjoint Representation
Given an element of the Lie algebra $X\in\mathfrak{g}$ and its corresponding Lie group $\mathcal{G}$, every $a\in\mathcal{G}$ defines an automorphism of the Lie algebra $Ad_{a}:\mathfrak{g}\rightarrow\mathfrak{g}$ by $Ad_{a}(X)=aXa^{-1}$. This is called the adjoint action of the group $\mathcal{G}$ on the Lie algebra $\mathfrak{g}$. It represents the change of basis operations on the algebra. Since the adjoint $Ad_{a}$ is linear, we can find a matrix that maps the $\mathbb{R}^{m}$ representation of the Lie algebra to another. That is, for every $Ad_{a}$ and $X\in\mathfrak{g}$, we have
@@
with $Adm_{a}\in\mathbb{R}^{m\times m}$, $x=X^{\vee}$ and $Adm_{a}x=(aXa^{-1})^{\vee}$. This is an important property as it allows us to model the group adjoint action using a matrix multiplication on $\mathbb{R}^{m}$, which enables the adjoint equivariant layer design in Section 4.
This work focuses on the matrix Lie group, where the Lie bracket is defined by the commutator: $\left[X,Y\right]=XY-YX$. It is worth noticing that the Lie bracket is equivariant under the group adjoint action, i.e., $\left[Ad_{a}(X),Ad_{a}(Y)\right]=Ad_{a}(\left[X,Y\right]),\forall a\in\mathcal{G}$.
@@ 3.3 Killing Form
If a Lie algebra $\mathfrak{g}$ is of finite dimension and associated with a field $\mathbb{R}$, a symmetric bilinear form called the Killing form is defined as:
$B(X,Y):\mathfrak{g}\times\mathfrak{g}\rightarrow\mathbb{R},\quad(X,Y)\mapsto\operatorname{tr}(ad_{X}\circ ad_{Y})$ (6)
@@ 4 Methodology
The Lie Neurons generalize VN with 3-channel geometric dimension to networks with $K$-channel geometric dimension, where $K$ is the dimension any semisimple Lie algebra, and the networks are equivariant to the adjoint action of the corresponding Lie group. Similar to how VN takes 3-dimensional points as input, our networks take elements of the Lie algebra as input. While a Lie algebra $\mathfrak{g}$ is a vector space with non-trivial structures, $X\in\mathfrak{g}$ can be expressed as a $K$-dimensional vector $x=X^{\vee}\in\mathbb{R}^{K}$ with appropriate bases using (2). This representation is the core concept for Lie Neurons.
Lie Neurons operate on the $\mathbb{R}^{K}$ representation of the Lie algebra. We denote the input as $\mathbf{x}\in\mathbb{R}^{K\times C}$, where $C$ is the feature dimension. This can be viewed as $C$ Lie algebra elements in the $\mathbb{R}^{K}$ form. That is, the Lie Neurons model $f:\mathfrak{g}^{C_{1}}\rightarrow\mathfrak{g}^{C_{2}}$ can be instantiated as
@@
A Lie Neurons network can be constructed with linear layers, two types of nonlinear activation layers, geometric channel mixing layers, pooling layers, and invariant layers. We start by discussing the linear layers as follows.
@@ 4.2.1 LN-ReLU: Nonlinearity Based on the Killing Form
We can use an invariant function to construct an equivariant nonlinear layer. The VN leverages the inner product in a standard vector space, which is invariant to $\mathrm{SO(3)}$, to design a vector ReLU nonlinear layer. We generalize this idea by replacing the inner product with the negative of the Killing form. As described in Section 3, the negative Killing form falls back to the inner product for compact semisimple Lie groups, and it is invariant to the group adjoint action.
@@ 4.2.2 LN-Bracket: Nonlinearity Based on the Lie Bracket
We use two learnable weight matrices $U,V\in\mathbb{R}^{C\times C}$ to map the input to different Lie algebra vectors, $\mathbf{u}=\mathbf{x}U,\mathbf{v}=\mathbf{x}V$. The Lie bracket of $\mathbf{u}$ and $\mathbf{v}$ becomes a nonlinear function on the input: $\mathbf{x}\mapsto[(\mathbf{x}U)^{\wedge},(\mathbf{x}V)^{\wedge}]^{\vee},\mathbb{R}^{K\times C}\rightarrow\mathbb{R}^{K\times C}$. Theoretically, we can use this as our nonlinear layer. However, we note that the Lie bracket essentially captures the failure of matrices to commute (Guggenheimer, 2012), and that $\left[X,X\right]=0,\forall X$. We find that when using two learnable Lie algebra elements from the same input, the Lie bracket cancels out most of the information and only passes through the residual. As a result, we add a skip connection to enhance the information flow, inspired by ResNet (He et al., 2016). The final design of the LN-Bracket layer becomes:
@@ 4.6 Relationship to Vector Neurons
Our method can almost specialize to the Vector Neurons when working with $\mathfrak{so}(3)$. This is because the linear adjoint matrix $Adm_{a}$ is exactly the rotation matrix for $\mathfrak{so}(3)$. Therefore, the group adjoint action becomes a left multiplication on the $\mathbb{R}^{3}$ representation of the Lie algebra. Moreover, $\mathrm{SO}(3)$ is a compact group. Thus, the negative Killing form of $\mathfrak{so}(3)$ defines an inner product. However, we omit the normalization in VN’s ReLU layer because the norm is not well defined when the Killing form is not negative definite. We also do not have a counterpart to VN’s batch normalization layer for the same reason.
Despite the similarity in appearance, the $\mathbb{R}^{3}$ vectors are viewed as points in the 3D Euclidean space in Vector Neurons, while they are treated as $\mathfrak{so}(3)$ Lie algebras in our framework, enabling a novel Lie bracket nonlinear layer. The geometric channel mixing layer also makes the model more flexible, which can be applied in both LN and VN.
@@ 5 Experiments
Our framework applies to arbitrary semi-simple Lie algebras. We conduct experiments on $\mathfrak{so}(3)$ and $\mathfrak{sl}(3)$ to validate its general applicability and effectiveness in various tasks.

### v3

@@ Abstract
This paper proposes an equivariant neural network that takes data in any finite-dimensional semi-simple Lie algebra as input. The corresponding group acts on the Lie algebra as adjoint operations, making our proposed network adjoint-equivariant. Our framework generalizes the Vector Neurons, a simple $\mathrm{SO}(3)$-equivariant network, from 3-D Euclidean space to Lie algebra spaces, building upon the invariance property of the Killing form. Furthermore, we propose novel Lie bracket layers and geometric channel mixing layers that extend the modeling capacity. Experiments are conducted for the $\mathfrak{so}(3)$, $\mathfrak{sl}(3)$, and $\mathfrak{sp}(4)$ Lie algebras on various tasks, including fitting equivariant and invariant functions, learning system dynamics, point cloud registration, and homography-based shape classification. Our proposed equivariant network shows wide applicability and competitive performance in various domains.
@@ 1 Introduction
The equivariance property preserves the symmetry group structure, often a Lie group, such that the feature map commutes with the group representation. An equivariant model, by construction, generalizes over the variations caused by the group actions. Therefore, it reduces the sampling complexity in learning and improves the robustness and transparency facing input variations. Equivariant models have gained success in various domains, including but not limited to the modeling of molecules (Thomas et al., 2018), physical systems (Finzi et al., 2020), social networks (Maron et al., 2018), images (Worrall et al., 2017), and point clouds (Zhu et al., 2023b). Convolutional neural networks (CNNs) are translation-equivariant, enabling stable image features regardless of the pixel positions in the image plane. Typical extensions include rotation (Cohen et al., 2017) and scale (Worrall & Welling, 2019) equivariance, while more general extensions are also explored (MacDonald et al., 2022).
Figure 1: Lie Neurons can be applied in learning dynamics. In this example, we learn the dynamics of a simulated free-rotation International Space Station, which we pose as an initial value problem in the Neural ODE framework. This figure shows the estimated trajectories and the learned vector field from the Lie Neurons. Detail descriptions can be found in Section 5.2.2.
In this paper, we propose a new type of equivariant model that captures the symmetry in Lie algebra spaces. For a given Lie algebra, we represent its elements as vectors by specifying a set of bases, and represent the adjoint actions as matrix multiplications, exploiting the isometry structure of Lie algebras to handle such data in typical vector form. Through the connection between inner products and the Killing form, we generalize the architecture of Vector Neurons (Deng et al., 2021), a $\mathrm{SO}(3)$-equivariant network originally designed for point cloud data in 3-D Euclidean space, to process data in arbitrary semisimple Lie algebra. We further build new types of equivariant network layers that exploit the structure of Lie algebras and extend the flexibility of the model. Using $\mathfrak{so}(3)$, $\mathfrak{sl}(3)$, and $\mathfrak{sp}(4)$ as examples, we conduct experiments on various tasks in physics, computer vision, and general function fitting, and show that our proposed network has wide applicability and competitive performance.
@@
    We propose a new adjoint-equivariant network architecture, enabling the processing of finite-dimensional semisimple Lie algebraic input data. Such models frequently appear in real-world geometric problems.
@@
    The software implementation is available at https://github.com/UMich-CURLY/LieNeurons.
@@ 2 Related Work
Equivariant networks enable the model output to change in a predicted way as the input goes through certain transformations. Current equivariant convolutional networks can be roughly categorized as regular group convolution and steerable group convolution. The regular group convolution exploits discretized subgroups to design equivariant convolutions. It is first introduced in Cohen & Welling (2016a), in which 90-degree discretizations of the $SO(2)$ are designed for 2D image processing. The approach is generalized to other discretized groups in $SE(2)$, $SE(3)$, $E(3)$, and $SIM(n)$ (Hoogeboom et al., 2018; Winkels & Cohen, 2018; Worrall & Brostow, 2018; Chen et al., 2021; Zhu et al., 2023a; Knigge et al., 2022). Steerable convolution is proposed in Cohen & Welling (2016b), leveraging the irreducible representations to remove the need for discretization and facilitate equivariant convolution on continuous groups in the frequency domain (Worrall et al., 2017; Cohen et al., 2017; Weiler et al., 2018; Thomas et al., 2018). Beyond convolutions, more general equivariant network architectures are proposed. For example, Fuchs et al. (2020); Hutchinson et al. (2021); Chatzipantazis et al. (2022) for transformers and Batzner et al. (2022); Brandstetter et al. (2021) for message passing networks. More recently, Geiger & Smidt (2022) introduces a generalized library for $E(3)$ equivariance based on variants of steerable equivariant networks. Vector Neurons (Deng et al., 2021) present a multi-layer perception (MLP) and graph network that generalize the scalar features to 3D features to realize $\mathrm{SO}(3)$-equivariance on spatial data. In addition to group-specific methods, more general recipes for building equivariant layers that are not limited to a specific group are also proposed (Kondor & Trivedi, 2018; Cohen et al., 2019; Weiler & Cesa, 2019; Xu et al., 2022; Lang & Weiler, 2020; Bekkers, 2019). The extension of equivariance beyond compact groups is also explored. Finzi et al. (2021) constructs MLPs equivariant to arbitrary matrix groups using their finite-dimensional representations. With the Monte Carlo estimator, equivariant convolutions are generalized to matrix groups with surjective exponential maps (Finzi et al., 2020) and all finite-dimensional Lie groups (MacDonald et al., 2022), where Lie algebras are used to parameterize elements in the continuous Lie groups as a lifted domain from the input space. More recently, Mironenco & Forré (2023) proposes a theoretic framework for $SL(n)$-equivariant convolution by decomposing the larger groups into manageable subgroups.
Our model structure resembles the MLP style of Vector Neurons (Deng et al., 2021), but our work models the equivariance of arbitrary semisimple groups under adjoint actions. Different from Vector Neurons, the input domain of our method is the Lie algebra. When working with $\mathfrak{so}(3)$, our method specializes to Vector Neurons, with an additional nonlinearity and a geometric mixing layer.
@@ 3 Preliminaries
We provide some preliminaries for Lie groups by focusing on matrix Lie groups. For detailed explanations, we refer the readers to Hall (2013); Rossmann (2006); Kirillov (2008).
@@ 3.1 Lie Group and Lie Algebra
A Lie group $\mathcal{G}$ is a smooth manifold whose elements satisfy the group axioms. The tangent space at the identity of a Lie group is named Lie algebra, denoted $\mathfrak{g}$. A Lie algebra locally captures the structure of the Lie group.
Every Lie algebra is equipped with an antisymmetric binary operator called the Lie bracket:
@@
In this work, we focus on finite-dimensional Lie algebras. Since Lie algebra is a vector space, one can always find a set of basis $E_{i}\in\mathfrak{g}$, which are linearly independent matrices of the vector space. Once we find the basis, we can represent each Lie algebra element as a linear combination of such basis by collecting the coefficients of each basis into a $\mathbb{R}^{m}$ vector (Chirikjian, 2011):
@@
where $E_{i}\in\mathfrak{g}$ are linear independent basis in $\mathfrak{g}$, and $e_{i}$ are the canonical basis of $\mathbb{R}^{m}$. An example of the Hat and Vee operation is provided in Appendix A.1.
@@ 3.2 Adjoint Representation
Given an element of the Lie algebra $X\in\mathfrak{g}$ and its corresponding Lie group $\mathcal{G}$, every $a\in\mathcal{G}$ defines an automorphism of the Lie algebra $Ad_{a}:\mathfrak{g}\rightarrow\mathfrak{g}$ by $Ad_{a}(X)=aXa^{-1}$. This is called the adjoint representation of the group $\mathcal{G}$ on the Lie algebra $\mathfrak{g}$. It amounts to the change of basis operations on the algebra. Since the adjoint $Ad_{a}$ is linear, we can find a matrix that maps the $\mathbb{R}^{m}$ form of the Lie algebra to another. That is, for every $Ad_{a}$ and $X\in\mathfrak{g}$, we have
@@
with $Adm_{a}\in\mathbb{R}^{m\times m}$, $x=X^{\vee}$ and $Adm_{a}x=(aXa^{-1})^{\vee}$. This is an important property as it allows us to model the group adjoint action using a matrix multiplication on $\mathbb{R}^{m}$, which enables the adjoint equivariant layer design.
Similarly, we can obtain the adjoint representation of the Lie algebra as $ad_{X}:\mathfrak{g}\rightarrow\mathfrak{g}$ by $ad_{X}(Y)=\left[X,Y\right]$. This work focuses on the matrix Lie group, where the Lie bracket is defined by the commutator: $\left[X,Y\right]=XY-YX$. It is worth noticing that the Lie bracket is equivariant under the group adjoint action, i.e., $\left[Ad_{a}(X),Ad_{a}(Y)\right]=Ad_{a}(\left[X,Y\right]),\forall a\in\mathcal{G}$.
@@ 3.3 Killing Form
If a Lie algebra $\mathfrak{g}$ is of finite dimension and associated with a field $\mathbb{R}$, a symmetric bilinear form called the Killing form is defined as (Kirillov, 2008):
$B(X,Y):\mathfrak{g}\times\mathfrak{g}\rightarrow\mathbb{R},\quad(X,Y)\mapsto\operatorname{tr}(ad_{X}\circ ad_{Y}).$ (5)
@@ 4 Methodology
Lie Neurons generalize VN with 3-channel geometric dimension to networks with $K$-channel geometric dimension, where $K$ is the dimension of any semisimple Lie algebra, and the networks are equivariant to the adjoint action of the corresponding Lie group. Similar to how VN takes 3-dimensional points as input, our networks take elements of the Lie algebra as input. While a Lie algebra $\mathfrak{g}$ is a vector space with non-trivial structures, $X\in\mathfrak{g}$ can be expressed as a $K$-dimensional vector $x=X^{\vee}\in\mathbb{R}^{K}$ with appropriate bases using (A.1). This $\mathbb{R}^{K}$ form is the core concept for Lie Neurons.
Lie Neurons operate on the $\mathbb{R}^{K}$ form of the Lie algebra. We denote the input as $\mathbf{x}\in\mathbb{R}^{K\times C}$, where $C$ is the feature dimension. This can be viewed as $C$ Lie algebra elements in the $\mathbb{R}^{K}$ form. That is, a Lie Neurons model $f:\mathfrak{g}^{C_{1}}\rightarrow\mathfrak{g}^{C_{2}}$ can be instantiated as
@@
A Lie Neuron network can be constructed with linear layers, two types of nonlinear activation layers, geometric channel mixing layers, pooling layers, and invariant layers. We start by discussing the linear layers as follows.
@@ 4.2.1 LN-ReLU: Nonlinearity Based on the Killing Form
We can use an invariant function to construct an equivariant nonlinear layer. The VN leverages the inner product in a standard vector space, which is invariant to $\mathrm{SO(3)}$, to design a vector ReLU nonlinear layer. The idea is to ”project” the negative part of a vector back to the zero plane, similar to how a 1-D ReLU rectifies the negative values. We generalize this idea by replacing the inner product with the negative of the Killing form. As described in Section 3, the negative Killing form falls back to the inner product for compact semisimple Lie groups, and it is invariant to the group adjoint action.
@@ 4.2.2 LN-Bracket: Nonlinearity Based on the Lie Bracket
We use two learnable weight matrices $U,V\in\mathbb{R}^{C\times C}$ to map the input to different Lie algebra vectors, $\mathbf{u}=\mathbf{x}U,\mathbf{v}=\mathbf{x}V$. The Lie bracket of $\mathbf{u}$ and $\mathbf{v}$ becomes a nonlinear function on the input: $\mathbf{x}\mapsto[(\mathbf{x}U)^{\wedge},(\mathbf{x}V)^{\wedge}]^{\vee},\mathbb{R}^{K\times C}\rightarrow\mathbb{R}^{K\times C}$. Theoretically, we can use this as our nonlinear layer. However, we note that the Lie bracket essentially captures the failure of matrices to commute (Guggenheimer, 2012), and that $\left[X,X\right]=0,\forall X$. Thus, the bracket captures the “non-commutativity” created by the two linear maps, which can be small in practice. As a result, we add a skip connection to enhance the information flow, inspired by ResNet (He et al., 2016). The final design of the LN-Bracket layer becomes:
@@ 4.6 Relationship to Vector Neurons
Our method can specialize to the Vector Neurons when working with $\mathfrak{so}(3)$. This is because the linear adjoint matrix $Adm_{a}$ is exactly the rotation matrix for $\mathfrak{so}(3)$. Therefore, the group adjoint action becomes a left multiplication on the $\mathbb{R}^{3}$ form of the Lie algebra. Moreover, $\mathrm{SO}(3)$ is a compact group. Thus, the negative Killing form of $\mathfrak{so}(3)$ defines an inner product. However, we omit the normalization in VN’s ReLU layer because the norm is not well defined when the Killing form is not negative definite. We also do not have a counterpart to VN’s batch normalization layer for the same reason.
Despite the similarity in appearance, the $\mathbb{R}^{3}$ vectors are viewed as points in the 3D Euclidean space in Vector Neurons, while they are treated as $\mathfrak{so}(3)$ Lie algebras in our framework, enabling a novel Lie bracket nonlinear layer. The geometric channel mixing layer also makes the model more flexible.
@@ 5 Experiments
Our framework applies to arbitrary semi-simple Lie algebras. We conduct experiments on $\mathfrak{so}(3)$ and $\mathfrak{sl}(3)$ to validate its general applicability and effectiveness in various tasks. Implementation details and additional experiments on $\mathfrak{sl}(3)$ can be found in the Appendix.
5.1 Experiment on $\mathfrak{sp}(4)$
We conduct an experiment on the Symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$. The symplectic Lie algebra can be defined as:
$\mathfrak{sp}(2n,\mathbb{R})=\{X\in GL(2n,\mathbb{R})\mid MX+X^{\intercal}M=0\},$ (18)
with $M=\begin{bmatrix}0&I_{n}\\ -I_{n}&0\end{bmatrix}$. The symplectic group and algebra can be used to model Hamilton mechanics and energy-conservation systems.
In this task, we aim to regress an invariant function
$g(X,Y)$ $=\sin(Tr(XY))+\cos(Tr(YY))$ (19)
$-\frac{Tr(YY)^{3}}{2}+\det(XY)+\exp(Tr(XX)),$
where $X,Y\in\mathfrak{sp}(4,\mathbb{R})$ and $Tr(\cdot)$ is the trace of the matrix. We generate 10,000 training and 10,000 testing data. In addition, during test time, we generate 500 $SP(4)$ matrices to perform test time augmentation. We report two MLP models with different feature dimensions for comparison. We also perform data augmentation during training for the MLP models. The invariance error captures the output consistency after the input is augmented.
From Table 1, we can see that Lie Neurons obtain superior accuracy while maintaining a low number of parameters compared to the MLP methods. In addition, the invariance error remains low without data augmentation, which confirms the invariant-by-construction property of Lie Neurons on the symplectic Lie algebra.
Table 1: The mean squared errors and the invariance errors in $\mathfrak{sp}(4)$ invariant function regression.
$Id$ $\mathrm{SP}(4)$
AVG $\downarrow$ AVG $\downarrow$ AVG $\downarrow$
MLP 256 Id 137,217 0.126 1.360 0.722
MLP 256 SP(4) 137,217 0.192 0.587 0.476
MLP 512 Id 536,577 0.107 0.906 0.585
MLP 512 SP(4) 536,577 0.123 0.446 0.374
Lie Neurons Id 263,170 $\mathbf{2.70\times 10^{-4}}$ $\mathbf{2.70\times 10^{-4}}$ $\mathbf{2.00\times 10^{-4}}$